OpenStudy (anonymous):

A snail travels at a rate of 2.45 feet per minute. How far will the snail travel in 8 minutes?

OpenStudy (campbell_st):

so if the use the formula \[Speed = \frac{distance}{time}\] then distance = Speed x time

OpenStudy (texaschic101):

or you can set it up as a proportion.. 2.45 / 1 = x / 8 --- 2.45 to 1 min = x to 8 min cross multiply x = (2.45)(8) x = 19.6 so he travels 19.6 feet in 8 minutes